# Properties of observable mixed inertial/gravito-inertial modes in $\gamma$ Doradus stars

### v1

@@ 6 Discussion and Conclusion
We have used the code TOP to study the resonances between inertial modes in the convective core and gravito-inertial modes in the radiative zone in $\gamma$ Dor stars. As Ouazzani et al. (2020), we obtained dips in the $\Delta P_{\mathrm{co}}$-$s$ relation around spin parameters of gravito-inertial modes approximately predicted by a simple analytical model. We computed those dips for different series of modes, different rotation rates and evolutionary stages across the main sequence. We quantified the dips using a Lorentzian function as suggested in Tokuno & Takata (2022), which allowed us to determine the evolution of their basic properties with the rotation and the evolutionary stage of the star. We then proposed an empirical model for the dips evolution.
@@
We also compared our numerical results with the model of Tokuno & Takata (2022) showing that it correctly predicts that the dip width is proportional to the rotation rate but does not account for the variation of the dip central spin parameter with rotation. We present a new model which accounts for this variation and which selects the significant dips by including the geometrical matching between the inertial mode and the gravito-inertial modes. The model is not fully predictive but it is useful to support and discuss the proposed empirical relations.
@@
Our empirical model gives access to the Brunt-Väisälä frequency jump at the core interface as well as the resonant inertial mode frequency for stars presenting a clear dip in their period spacing patterns. Such dips have been observed, as shown in Saio et al. (2021) and further research could be led in order to find dips in a larger number of $\gamma$ Dor stars in the TESS and Kepler data. This could be challenging, as modes in the dips may be harder to observe due to their higher inertia, their energy being more concentrated in their core. As our empirical model relies on only two different stellar models, it needs to be tested on more stellar models in order to estimate its accuracy. Such study is not trivial as computing and identifying series of gravito-inertial modes with 2D oscillation codes can be a tedious process. The resonance phenomenon between the core and the envelope may not be exclusive to $\gamma$ Dor stars and similar dips structures could arise from $\Delta P$ $P$ diagram of other kind of stars. Slow Pulsating B (SPB) stars are good candidates, even though no dips have been observed yet (Aerts & Mathis, 2023). A thorough investigation taking into account the expected spin parameters of the resonant inertial modes would be needed to establish if resonances occur in the frequency ranges observed in SPB stars.
@@
The main limitation of our present numerical study is related to the modelling of the rotation: we neglected the distortion induced by the centrifugal force and assumed a uniform rotation profile. For rotation rates observed in $\gamma$ Dor stars, we expect a weak impact of the centrifugal distortion on our results, since it mainly affects the external layers of the stars: the structure of convective cores and the surrounding regions where g modes have large amplitudes remains quasi-spherical (see, e.g., Ballot et al., 2010, 2012). However, the influence of the differential rotation should be carefully investigated in a further study, especially when the core spins at a rotation which is different from the radiative zone. That kind of differential rotation has already been investigated and identified through the study of dips observed in 16 $\gamma$ Dor star in Saio et al. (2021). Although their differential rotation is shown to be small to non-significant in most cases, the core of one of the 16 stars was found to rotate 20% faster than the surrounding radiative zone. Such differential rotation would cause the frequency of the resonant inertial mode to be misestimated as its estimate relies on the use of the inner radiative zone rotation rate obtained through the TAR. Nevertheless, it remains unclear how such differential rotation would affect the width of the dips and the estimate of the Brunt-Väisälä jump at the core interface. If $\beta$ remains constant, the width is only related to the radiative zone rotation rate rather than the core rotation rate. It implies that the estimate of the Brunt-Väisälä frequency jump through the dip width and the rotation rate of the inner radiative zone should remain correct despite the presence of such differential rotation.
@@
We neglected non-adiabatic effects in our analysis. This approximation is supported by non-adiabatic calculations of high-order g modes which have been performed to investigate the excitation mechanism of $\gamma$ Dor pulsations (Guzik et al., 2000; Dupret et al., 2005; Bouabid et al., 2013). They show that the adiabatic analysis provides precise enough oscillation frequencies (Dupret et al., 2005). The spatial distribution of g modes are affected by non-adiabatic effects, but those are only significant in the upper layers of the star where the thermal relaxation time is smaller or of the same order as the oscillation period (Dupret et al., 2002, 2005). We thus expect that deep in the star, at the interface of the convective core and the radiative zone, the mode geometry is not affected by non-adiabatic effects. Thus, the adiabatic approximation we used to model the coupling between modes in the convective core and in the radiative envelope should be precise enough.
@@
Our stellar structure models include turbulent diffusion to reproduce the different mixing processes that occur in stars. The value of diffusion coefficient we used is the one proposed by Ouazzani et al. (2020), based on a previous calibration to Genova models (Miglio et al., 2008). The mixing is thus very efficient and prevent the formation of sharp features in the Brunt-Väisälä frequency profile in the g-mode cavity. Such features are know to generate glitches, visible in $\Delta P$ $s$ diagrams as oscillations or periodic dips (see for example Miglio et al., 2008). While absent in our computations, glitches may be present in real data and perturb the characterisation or even the detection of dips. Conversely, dips caused by the resonance phenomenon can potentially be mistaken for a glitch signature and could lead to wrong estimations of stellar parameters (see Mombarg et al., 2021).
@@
The frequency and the width of the dips depends on the structure of the convective core through the spin parameter and the spatial distribution of the resonant core inertial modes. As shown by Eq. (23), it is the density stratification of the convective core that controls the inertial mode characteristics. Thus the evolutionary stage or the stellar mass affect the dips as far as they affect the convective core density stratification. It is worth noting that the core size in itself has no direct impact on the inertial modes. Further investigations are needed to understand the relation between the frequencies of the inertial modes in the core and the stellar properties. It would thus allow us to use the said frequencies extracted by our empirical model to constrain stellar parameters. In particular, coupled with the jump in the Brunt-Väisälä frequency at the core interface we extract from our model and the buoyancy radius deduced from the TAR, it could be a powerful tool to determine stellar ages. Determining precise and accurate ages will be crucial for the PLATO mission (Rauer et al., 2014), which will use $\gamma$ Dor stars as scientific calibrators.
@@ Acknowledgements.
We thank R.M. Ouazzani, M. Takata and D.R. Reese for very useful discussions and for performing some tests and comparisons, which were crucial to validate our computations with TOP. We acknowledge support from the Centre National d’Etudes Spatiales (CNES). This work was supported by the ”Programme National de Physique Stellaire” (PNPS) of CNRS/INSU co-funded by CEA and CNES.
@@ Appendix A Uniform density model
where $\mu=1/s=\omega_{\mathrm{co}}/2\Omega$ and $\ell_{i}$ is the degree of the inertial mode. The spin parameters of inertial modes that could couple with gravito-inertial series studied here are presented in Table 7. For an inertial mode in the core to be significantly coupled with a gravito-inertial mode in the radiative zone, there needs to be spatial and temporal correspondences. The two coupling modes must share a similar latitudinal profile at the interface between the core and the radiative zone and must oscillate with a similar frequency. The condition on the frequency is easily fulfilled as the frequency spectrum of gravito-inertial modes is dense. To quantify the spatial correspondence between an inertial and a gravito-inertial mode, Ouazzani et al. (2020) uses a correlation coefficient between the associated Legendre polynomial (which describes the inertial mode in the core in the analytical model) and the Hough function (which describes the gravito-inertial mode in the envelope in the TAR approximation). Those coefficient are presented in parenthesis next to the spin parameter of the modes in Table 7. The modes $(\ell_{i}=3$ $m=-1$), $(\ell_{i}=3$ $m=0$) and $(\ell_{i}=4$ $m=-2$) with the highest correlation coefficients are shown in Table 3.
Table 7: Spin parameters of inertial modes obtained using an analytical uniform density core model. The correlation coefficient between the inertial mode and the Hough function of the same symmetry class and spin parameter is shown in parenthesis.
@@ Appendix B Core-only model
We run TOP with a truncated version of our models to obtain the resonant frequencies. Table 8 shows the spin parameters of the inertial modes coupling with the series of gravito-inertial modes ($\ell=1$ $m=-1$), ($\ell=1$ $m=0$) and ($\ell=2$ $m=-2$) for the 3 models. They are labelled as ($\ell_{i}=3$ $m=-1$), ($\ell_{i}=3$ $m=0$) and ($\ell_{i}=4$ $m=-2$) respectively as they remain similar to the modes calculated with the uniform density model presented in Appendix A (see Table 3). Despite their apparent resemblance, it has to be noted that their spin parameters vary significantly with the density stratification as the relative differences between the 2m and 3t models’ spin parameters and the uniform-density model exceed 20%. To stress their differences, we show the mode profiles along the equator for both the uniform-density and 3t models in Fig. 18.
@@
Table 8: Spin parameters $s_{*}$ of the resonant inertial modes coupling with the series of modes ($\ell=1$ $m=-1$), ($\ell=1$ $m=0$) and ($\ell=2$ $m=-2$), for the models 1z, 2m and 3t using a core-only model at $\Omega=0.1\Omega_{K}$ and $\Omega=0.5\Omega_{K}$.
@@
Table 9: Inertial modes in the core coupling with the ($\ell=1$ $m=-1$), ($\ell=1$ $m=0$) and ($\ell=2$ $m=-2$) series of gravito-inertial modes in the envelope for the uniform density and 3t models.
@@
Figure 18: Eulerian pressure perturbations normalised by their value at the core radius $R_{\mathrm{c}}$ of the ($\ell_{i}=3$, $m=-1$) mode, as a function of the radius at the colatitude $\theta=\pi/2$ (equator), for the uniform-density model (blue) and for the 3t model (red).
@@ Appendix C Results and analysis for the ($\ell=1$, $m=0$) and ($\ell=2$, $m=-2$) series
The dips in the series of modes ($\ell=1$, $m=0$) computed for the models 1z, 2m and 3t at different rotations are presented in Figs. 19, 20 and 21.
@@
The evolution of the width $\sigma$ and location $s_{c}$ of dips as a function of the rotation are shown in Figs. 22 and 23. In the latter, we also show the evolution of the spin parameter $s_{*}$ of the resonant inertial mode of core-only models and the spin parameter of the same mode in a uniform-density sphere.
@@
Figure 25: Ratio $s_{c}/s_{*}$ as a function of $\Omega/N_{0}$ ($\epsilon$) for the series ($\ell=1$ $m=0$), for the 1z, 2m and 3t models. The purple dashed line shows the fit presented in Table 4 for the three models together and the purple area shows the related error. The x-axis is logarithmic for visualisation purposes.
@@
We used the empirical relations (10) and (12) on our numerical calculations to estimate $s_{*}$ and $N_{0}$ and compared them to their true values. The errors made are shown in Figs. 27 and 28. In the first we used the three models, in the latter only the models 1z and 2m. Relative errors behave quite differently when compared to the $\ell=|m|$ (Kelvin) modes. The estimate for $N_{0}$ is bad for the 3t model regardless of the rotation rate, making the estimate of $s_{*}$ also worse for the 3t model.
Figure 27: Relative errors for the series ($\ell=1$ $m=0$) on the estimated $N_{0}^{\mathrm{calc}}$ (upper panel) and $s_{*}^{\mathrm{calc}}$ (lower panel) using the three models 1z, 2m and 3t.
@@
The dips in the series of modes ($\ell=2$, $m=-2$) computed for the models 1z, 2m and 3t at different rotations are presented in Figs. 29, 30 and 31.
@@
The evolution of the width $\sigma$ and location $s_{c}$ of dips as a function of the rotation are shown in Figs. 32 and 33. In the latter, we also show the evolution of the spin parameter $s_{*}$ of the resonant inertial mode of core-only models and the spin parameter of the same mode in a uniform-density sphere.
@@
In Fig. 36, we compare the relative evolution of $\sigma$ with rotation in our numerical computations and in the theoretical model of Tokuno & Takata (2022). As for the series ($\ell=1$, $m=-1$, while $\sigma$ is clearly linear for the 1z and 2m models, it is not the case for the 3t model. We can see that this tendency is more pronounced for the ($\ell=2$, $m=-2$) series than it is for the ($\ell=1$, $m=-1$) series.
@@
We used the empirical relations (10) and (12) on our numerical calculations to estimate $s_{*}$ and $N_{0}$ and compared them to their true values. The errors made are shown in Figs. 37 and 38. In the first we used the three models, in the latter only the models 1z and 2m. We can see that the relative errors on $N_{0}$ and $s_{*}$ follow the same kind of trend as for the ($\ell=1$, $m=-1$) series.
@@
We present an analytical model of mixed inertial/gravito-inertial modes when the density is assumed to be uniform in the convective core. This model is similar to the one proposed by Tokuno & Takata (2022) in that its uses the same hypothesis to derive analytical solutions in the convective core and in the radiative zone respectively. But it goes a step further by improving the continuity of the mixed modes at the convective/radiative interface.

### v2

@@ 6 Discussion and conclusion
We have used the code TOP to study the resonances between inertial modes in the convective core and gravito-inertial modes in the radiative zone in $\gamma$ Dor stars. As in Ouazzani et al. (2020), we obtained dips in the $\Delta P_{\mathrm{co}}-s$ relation around spin parameters of gravito-inertial modes approximately predicted by a simple analytical model. We computed those dips for different series of modes and different rotation rates and evolutionary stages across the main sequence. We quantified the dips using a Lorentzian function as suggested in Tokuno & Takata (2022), which allowed us to determine the evolution of their basic properties with the rotation and the evolutionary stage of the star. We then proposed an empirical model for the evolution of the dips.
@@
We also compared our numerical results with the model of Tokuno & Takata (2022), showing that it correctly predicts that the dip width is proportional to the rotation rate, but it does not account for the variation of the dip central spin parameter with rotation. We have presented a new model that accounts for this variation and selects the significant dips by including the geometrical matching between the inertial mode and the gravito-inertial modes. The model is not fully predictive, but it is useful to support and discuss the proposed empirical relations.
@@
Our empirical model gives access to the Brunt-Väisälä frequency jump at the core interface as well as the resonant inertial mode frequency for stars presenting a clear dip in their period spacing patterns. Such dips have been observed, as shown in Saio et al. (2021), and further research could be led in order to find dips in a larger number of $\gamma$ Dor stars in the TESS and Kepler data. This could be challenging, as modes in the dips may be harder to observe due to their higher inertia, as their energy is more concentrated in their core. As our empirical model relies only on two different stellar models, it needs to be tested on more stellar models in order to estimate its accuracy. Such a study is not trivial, as computing and identifying series of gravito-inertial modes with 2D oscillation codes can be a tedious process. The resonance phenomenon between the core and the envelope may not be exclusive to $\gamma$ Dor stars, and similar dip structures could arise from the $\Delta P$ $P$ diagram of other types of stars. Slow pulsating B (SPB) stars are good candidates even though no dips have been observed yet (Aerts & Mathis, 2023). A thorough investigation taking into account the expected spin parameters of the resonant inertial modes would be needed to establish whether resonances occur in the frequency ranges observed in SPB stars.
@@
The main limitation of our present numerical study is related to the modelling of the rotation: We neglected the distortion induced by the centrifugal force and assumed a uniform rotation profile. For rotation rates observed in $\gamma$ Dor stars, we expect a weak impact of the centrifugal distortion on our results since it mainly affects the external layers of the stars. Indeed, the structure of convective cores and the surrounding regions where g modes have large amplitudes remains quasi-spherical (see, e.g., Ballot et al., 2010, 2012). However, the influence of the differential rotation should be carefully investigated in a further study, especially when the core spins at a rotation that is different from the radiative zone. That kind of differential rotation has already been investigated and identified through the study of dips observed in 16 $\gamma$ Dor stars in Saio et al. (2021). Although their differential rotation is shown to be small to non-significant in most cases, the core of one of the 16 stars was found to rotate 20% faster than the surrounding radiative zone. Such differential rotation would cause the frequency of the resonant inertial mode to be misestimated, as its estimate relies on the use of the inner radiative zone rotation rate obtained through the TAR. Nevertheless, how such differential rotation would affect the width of the dips and the estimate of the Brunt-Väisälä jump at the core interface remains unclear. If $\beta$ remains constant, the width is only related to the radiative zone rotation rate rather than the core rotation rate. It implies that the estimate of the Brunt-Väisälä frequency jump through the dip width and the rotation rate of the inner radiative zone should remain correct despite the presence of such differential rotation.
@@
We neglected non-adiabatic effects in our analysis. This approximation is supported by non-adiabatic calculations of high-order g modes, which have been performed to investigate the excitation mechanism of $\gamma$ Dor pulsations (Guzik et al., 2000; Dupret et al., 2005; Bouabid et al., 2013). These calculations showed that the adiabatic analysis provides precise enough oscillation frequencies (Dupret et al., 2005). The spatial distribution of the g modes is affected by non-adiabatic effects, but those are only significant in the upper layers of the star, where the thermal relaxation time is smaller or of the same order as the oscillation period (Dupret et al., 2002, 2005). We thus expect that deep in the star, at the interface of the convective core and the radiative zone, the mode geometry is not affected by non-adiabatic effects. Thus, the adiabatic approximation we used to model the coupling between modes in the convective core and in the radiative envelope should be precise enough.
@@
Our stellar structure models include turbulent diffusion to reproduce the different mixing processes that occur in stars. The value of the diffusion coefficient we used is the one proposed by Ouazzani et al. (2020), which is based on a previous calibration to Genova models (Miglio et al., 2008). The mixing is thus very efficient and prevents the formation of sharp features in the Brunt-Väisälä frequency profile in the g-mode cavity. Such features are known to generate glitches, visible in $\Delta P-s$ diagrams as oscillations or periodic dips (see for example Miglio et al., 2008). While absent in our computations, glitches may be present in real data and perturb the characterisation or even the detection of dips. Conversely, dips caused by the resonance phenomenon can potentially be mistaken for a glitch signature and could lead to wrong estimations of stellar parameters (see Mombarg et al., 2021).
@@
The frequency and the width of the dips depends on the structure of the convective core through the spin parameter and the spatial distribution of the resonant core inertial modes. As shown by Eq. (23), it is the density stratification of the convective core that controls the inertial mode characteristics. Thus, the evolutionary stage or the stellar mass affects the dips as far as they affect the convective core density stratification. It is worth noting that the core size itself has no direct impact on the inertial modes. Further investigations are needed to understand the relation between the frequencies of the inertial modes in the core and the stellar properties. Such studies would thus allow one to use the said frequencies extracted by our empirical model to constrain stellar parameters. In particular, coupled with the jump in the Brunt-Väisälä frequency at the core interface we extracted from our model and the buoyancy radius deduced from the TAR, the inertial mode frequencies could be a powerful tool to determine stellar ages. Determining precise and accurate ages will be crucial for the PLATO mission (Rauer et al., 2014), which will use $\gamma$ Dor stars as scientific calibrators.
@@ Acknowledgements.
We thank the anonymous referee for carefully reading our manuscript and providing helpful and constructive comments. We thank R.M. Ouazzani, M. Takata and D.R. Reese for very useful discussions and for performing some tests and comparisons, which were crucial to validate our computations with TOP. We acknowledge support from the Centre National d’Etudes Spatiales (CNES). This work was supported by the ”Programme National de Physique Stellaire” (PNPS) of CNRS/INSU co-funded by CEA and CNES.
@@ Appendix A Uniform density model
where $\mu=1/s=\omega_{\mathrm{co}}/2\Omega$ and $\ell_{i}$ is the degree of the inertial mode. The spin parameters of inertial modes that could couple with gravito-inertial series studied here are presented in Table 7. For an inertial mode in the core to be significantly coupled with a gravito-inertial mode in the radiative zone, there needs to be spatial and temporal correspondences. The two coupling modes must share a similar latitudinal profile at the interface between the core and the radiative zone and must oscillate with a similar frequency. The condition on the frequency is easily fulfilled as the frequency spectrum of gravito-inertial modes is dense. To quantify the spatial correspondence between an inertial and a gravito-inertial mode, Ouazzani et al. (2020) uses a correlation coefficient between the associated Legendre polynomial (which describes the inertial mode in the core in the analytical model) and the Hough function (which describes the gravito-inertial mode in the envelope in the TAR approximation). Those coefficient are presented in parentheses next to the spin parameter of the modes in Table 7. The modes $(\ell_{i}=3$ $m=-1$), $(\ell_{i}=3$ $m=0$), and $(\ell_{i}=4$ $m=-2$) with the highest correlation coefficients are shown in Table 7.
Table 7: Spin parameters of inertial modes obtained using an analytical uniform density core model. The correlation coefficient between the inertial mode and the Hough function of the same symmetry class and spin parameter is shown in parentheses.
@@ Appendix B Core-only model
We run TOP with a truncated version of our models to obtain the resonant frequencies. Table 8 shows the spin parameters of the inertial modes coupling with the series of gravito-inertial modes ($\ell=1$ $m=-1$), ($\ell=1$ $m=0$), and ($\ell=2$ $m=-2$) for the 3 models. They are labelled as ($\ell_{i}=3$ $m=-1$), ($\ell_{i}=3$ $m=0$), and ($\ell_{i}=4$ $m=-2$) respectively as they remain similar to the modes calculated with the uniform density model presented in Appendix A (see Table 7). Despite their apparent resemblance, it has to be noted that their spin parameters vary significantly with the density stratification as the relative differences between the 2m and 3t models’ spin parameters and the uniform-density model exceed 20%. To stress their differences, we show the mode profiles along the equator for both the uniform-density and 3t models in Fig. 18.
@@
Table 8: Spin parameters $s_{*}$ of the resonant inertial modes coupling with the series of modes ($\ell=1$ $m=-1$), ($\ell=1$ $m=0$), and ($\ell=2$ $m=-2$), for the models 1z, 2m, and 3t using a core-only model at $\Omega=0.1\Omega_{K}$ and $\Omega=0.5\Omega_{K}$.
@@
Table 9: Inertial modes in the core coupling with the ($\ell=1$ $m=-1$), ($\ell=1$ $m=0$), and ($\ell=2$ $m=-2$) series of gravito-inertial modes in the envelope for the uniform density and 3t models.
@@
Figure 18: Eulerian pressure perturbations normalised by their value at the core radius $R_{\mathrm{c}}$ of the ($\ell_{i}=3$, $m=-1$) mode, as a function of the radius at the colatitude $\theta=\pi/2$ (equator), for the uniform density model (blue) and for the 3t model (red).
@@ Appendix C Results and analysis for the ($\ell=1$, $m=0$) and ($\ell=2$, $m=-2$) series
The dips in the series of modes ($\ell=1$, $m=0$) computed for the models 1z, 2m, and 3t at different rotations are presented in Figs. 19, 20, and 21.
@@
The evolution of the width $\sigma$ and location $s_{c}$ of dips as a function of the rotation are shown in Figs. 22 and 23. In the latter, we also show the evolution of the spin parameter $s_{*}$ of the resonant inertial mode of core-only models and the spin parameter of the same mode in a uniform density sphere.
@@
Figure 25: Ratio $s_{c}/s_{*}$ as a function of $\Omega/N_{0}$ ($\epsilon$) for the series ($\ell=1$ $m=0$), for the 1z, 2m, and 3t models. The purple dashed line shows the fit presented in Table 4 for the three models together and the purple area shows the related error. The x-axis is logarithmic for visualisation purposes.
@@
We used the empirical relations (10) and (12) on our numerical calculations to estimate $s_{*}$ and $N_{0}$ and compared them to their true values. The errors made are shown in Figs. 27 and 28. In the first one, we used the three models and in the latter only the models 1z and 2m. Relative errors behave quite differently when compared to the $\ell=|m|$ (Kelvin) modes. The estimate for $N_{0}$ is bad for the 3t model regardless of the rotation rate, making the estimate of $s_{*}$ also worse for the 3t model.
Figure 27: Relative errors for the series ($\ell=1$ $m=0$) on the estimated $N_{0}^{\mathrm{calc}}$ (upper panel) and $s_{*}^{\mathrm{calc}}$ (lower panel) using the three models 1z, 2m, and 3t.
@@
The dips in the series of modes ($\ell=2$, $m=-2$) computed for the models 1z, 2m, and 3t at different rotations are presented in Figs. 29, 30 and 31.
@@
The evolution of the width $\sigma$ and location $s_{c}$ of dips as a function of the rotation are shown in Figs. 32 and 33. In the latter, we also show the evolution of the spin parameter $s_{*}$ of the resonant inertial mode of core-only models and the spin parameter of the same mode in a uniform density sphere.
@@
In Fig. 36, we compare the relative evolution of $\sigma$ with rotation in our numerical computations and in the theoretical model of Tokuno & Takata (2022). As for the series ($\ell=1$, $m=-1$, while $\sigma$ is clearly linear for the 1z and 2m models, it is not the case for the 3t model. One can see that this tendency is more pronounced for the ($\ell=2$, $m=-2$) series than it is for the ($\ell=1$, $m=-1$) series.
@@
We used the empirical relations (10) and (12) on our numerical calculations to estimate $s_{*}$ and $N_{0}$ and compared them to their true values. The errors made are shown in Figs. 37 and 38. In the first one, we used the three models and in the latter only the models 1z and 2m. One can see that the relative errors on $N_{0}$ and $s_{*}$ follow the same kind of trend as for the ($\ell=1$, $m=-1$) series.
@@
We present an analytical model of mixed (inertial and gravito-inertial modes) when the density is assumed to be uniform in the convective core. This model is similar to the one proposed by Tokuno & Takata (2022) in that its uses the same hypothesis to derive analytical solutions in the convective core and in the radiative zone respectively. However, it goes a step further by improving the continuity of the mixed modes at the convective-radiative interface.